# Kirchhoff Meets Johnson: In Pursuit of Unconditionally Secure Communication

### v2

@@ III-A Fundamentals of KLJN
One can ask the following question regarding the KLJN system: Compared to the bulletproof security provided by quantum weirdness, can a simple wire that is freely exposed to the public, with measurable voltage and current waveforms, offer unconditional security for eavesdroppers with unlimited measurement capabilities? The answer is yes, thanks to the laws of physics (thermodynamics). In other words, like its quantum-based counterpart, violating the security of KLJN requires violating specific laws of physics. One of the reasons why our community has not fully understood the scheme of KLJN in the past 15 years is due to its multidisciplinary nature, which includes statistical physics, circuit noise theory, stochastic signals, and information-theoretic security. In the following subsection, we will discuss the basic technical operation principles of the KLJN scheme.
@@ III-B KLJN Operation
which can alternatively be expressed in units of dBm/Hz or watts/Hz considering a reference impedance. Here, $R_{A}$ and $R_{B}$ represent the randomly selected resistors by Alice and Bob, respectively, that is, $R_{A},R_{B}\in\left\{R_{L},R_{H}\right\}$. Accordingly, for the four possible cases of selected Alice/Bob bits, namely $00,01,10,$ and $11$, where the first and second bits respectively stand for the selected bits by Alice and Bob, voltage waveform samples on the line will be Gaussian distributed random variables with different variances. The corresponding variance values of the noise samples are given as follows
@@
$\sigma_{11}^{2}$ $=4kT\frac{R_{H}R_{H}}{R_{H}+R_{H}}\Delta f=4kTR_{L}\frac{\alpha}{2}\Delta f,$
where $\alpha$ is the ratio of resistances, that is, $R_{H}=\alpha R_{L}$. As illustrated in Fig. 1, the fluctuations on the line will depend on the selected resistance values at both sides of the system. Specifically, the fluctuations on the line will be high if both Alice and Both select the large resistance $R_{H}$. On the other hand, if both Alice and Bob select the small resistance $R_{L}$, the fluctuations will be small. What about the intermediate level? If one user (Alice or Bob) selects the large resistance while the other selects the small one, the noise variance on the line takes an intermediate value, which lies at the heart of the KLJN operation [10]. This intermediate noise variance value creates an indeterminacy for an eavesdropper (Eve), thus paving the way for unconditional security. Before providing more detailed discussions on this important topic, let us give a numerical example of the noise variances of different cases. Specifically, for the value of $\alpha=10$, noise variances on the line for the cases of intermediate and high fluctuations would be $1.8182$ and $10$ times greater than that of small fluctuations, respectively. An unconditionally secure bit exchange occurs when the bit values (or equivalently resistances) at the two ends of the KLJN system differ [^1] Although a random key is established between Alice and Bob by the KLJN system, it has been widely regarded as a key exchanger in literature due to its unique architecture.. For this case, an intermediate level is observed for the noise voltage variance (mean-square noise voltage), as shown in Fig. 1. Even though this intermediate level can be easily detected by Eve from a sufficient number of noise samples taken over the line, the specific contributions of Alice and Bob cannot be understood. This confusion on Alice’s and Bob’s resistances, created deliberately for Eve, ensures unconditional security at the level of quantum-based key distribution systems. In other words, the KLJN scheme consistently creates an indeterminacy for external observers, rendering the intermediate noise variance meaningless. In contrast, the cases of $00$ and $11$, respectively, provide the lowest and highest noise variances on the link. In this case, an eavesdropper can identify Alice/Bob bits from its measurements without a significant effort, corresponding to a non-secure bit exchange. As stated earlier, under ideal conditions, the cases of $01$ and $10$ produce equivalent noise variances for the noise samples. As a result, the specific locations of Alice/Bob bits cannot be determined from measurements taken on the line between the users. This unique property of KLJN provides a sort of powerful encryption and ensures absolute (unbreakable) security. For Alice and Bob, determining their partner’s selected bit is straightforward for all four cases using voltage and/or current samples and the knowledge of their own bit. However, they have to take a sufficient number of noise samples for each transmission interval to reduce their bit errors. While we provide brief coverage of this issue in the following subsection, interested readers are referred to [3] for a comprehensive theoretical evaluation of KLJN’s bit error probability (BEP) and its different types of detectors exploiting noise voltage and/or current samples. It is worth noting that the switching mechanism of the KLJN scheme among low and high resistances has conceptual similarities with the index modulation (IM) method [11], which embeds information in certain system entities such as antennas, subcarriers, and time slots. To be more specific, a KLJN communicator performs a unique form of IM by considering the indices of its available two resistors, or equivalently, for three possible noise voltage (or current) power spectral densities; however, it requires a no-wave solution for security. As in all IM systems, index selection does not consume power, and the primary task of the users is to extract the information embedded in the corresponding system entity through IM, which is the noise variance for the case of the KLJN scheme. In this context, selecting the number of noise samples per bit is critical, as discussed in the following subsection. While concluding this subsection here, Section VI discusses the state-of-the-art KLJN solutions and their security vulnerabilities.
@@ III-C Conquering the Noise: KLJN Detection
Noise: an enemy to be dealt with for achieving reliable and secure communication. Deviating from mainstream engineering, we can ask the following question: is it possible to use the noise itself, for instance, as biology does using noise-alike signals as the information carrier, for communication purposes? KLJN uses the noise itself for this purpose; as a result, we need to explore practical ways to extract the information embedded in this noise. In this subsection, we shed light on this issue. To determine their partner’s bit in the KLJN system, Alice and Bob take samples from the thermal noise voltage (or current) on the wire for each bit duration $(T_{b})$. The total number of “uncorrelated” noise samples taken per bit, denoted by $N$, can be derived from the Wiener–Khinchin theorem, considering the sinc-type time autocorrelation function of the band-limited white noise. Accordingly, for a noise bandwidth of $\Delta f$ Hz, a maximum of $N=2T_{b}\Delta f$ noise samples can be taken for each bit duration to ensure that the samples are statistically independent, which is also consistent with the Nyquist’s sampling theorem. In other words, in accordance with the sampling theorem, faster-than-Nyquist sampling would result in correlated noise samples. Consequently, to increase $N$ to improve the accuracy of noise variance estimates, one needs to increase $T_{b}$; however, this reduces the bit rate given by $R_{b}=1/T_{b}$. This provides an interesting trade-off between the bit error rate (BER) and data rate for the KLJN system. Another attractive feature of KLJN is that no matter how strong Eve’s signal processing capabilities are, the same limit also applies to her; she cannot gain further knowledge by expanding her observation interval or increasing the number of noise samples.
@@
$\hat{\sigma}^{2}=\frac{1}{N}\sum\nolimits_{k=1}^{N}x_{k}^{2}.$
An effective way for Alice and Bob to detect their partner’s bit is to employ a threshold-based detector, as shown in Fig. 2, by employing two threshold values, $\gamma_{1}$ and $\gamma_{2}$. At this point, considering the value of $\hat{\sigma}^{2}$ as well as their own bit, Alice and Bob can determine their partner’s bit. For instance, assume that Alice’s bit is 00. In this case, she considers the first threshold, $\gamma_{1}$, to detect Bob’s bit. If $\hat{\sigma}^{2}<\gamma$, her decision becomes 00, and she marks the corresponding bit as an insecure one since she assumes that the selected bit combination is $00$. Particularly, she is aware of the fact that a third party can easily identify the case of $00$ through measurements on the line. On the contrary, if $\hat{\sigma}^{2}>\gamma$, her decision becomes $1$, which corresponds to an unconditionally secure bit exchange (generation). Since Alice and Bob’s bits are different for this secure case, one can flip his/her bits to create a shared and secure key after the transmission. It is worth noting that for the detection scheme of Fig. 2, the error events between $00$ and $01/10$ dominate the overall BEP due to the proximity of the corresponding variance values on the decision space. Employing joint or alternating noise current samples, which provide an opposite relationship between noise variances compared to those given by voltage samples, this problem can be alleviated. As a result, much lower BEP can be achieved using more sophisticated detectors. It can be shown that the BEP of the KLJN scheme exponentially decays with the number of noise samples per bit, that is, $P_{b}\propto\exp(-N)$, which is promising to obtain very low BEP values for secure key generation between two partners. Furthermore, lowering the BEP does not compromise the unconditional security of the system. A detailed theoretical framework is represented in [3] by considering decision errors due to the randomness of the estimated noise variance. Furthermore, two new detectors are also presented in this study by exploiting joint voltage and noise samples further to improve the BEP performance of the KLJN scheme.
@@ IV In Pursuit of Unconditional Security
One of the objectives during the introduction of KLJN in the early 2000s had been to challenge the following status quo on unconditional security: Only quantum physics can provide unconditional security for the key distribution. At this point, the conceptual difference between post-quantum cryptography and quantum cryptography is worth noting. Specifically, post-quantum cryptography, also known as quantum-resistant or quantum-safe cryptography, deals with quantum-resistant cryptographic algorithms, which are implemented on existing platforms and derive their security through mathematical complexity. In principle, these algorithms will be resilient to attacks from quantum computers thanks to their highly complicated mathematical frameworks [12]. On the other hand, quantum cryptography, specifically quantum key distribution (QKD), uses principles of quantum mechanics to distribute the keys to be used later in a suitable form of cryptography and is genuinely unbreakable, unlike mathematical encryption. In other words, the security level provided by QKD is beyond the attackers of the future who possess quantum computers and even have infinite computational power [^2] This strong assertion holds when we employ one-time pad ciphers with keys established using QKD, which will be discussed subsequently.. In what follows, we dive into a historical journey on secure communication to catch valuable insights on the motivation of these physics-based key generation systems, KLJN and QKD, and their limitations in practice. The desire to communicate secretly has existed as long as writing and encrypted communication methods were developed in many ancient societies [13]. Notable historical examples include the scytale of Spartans from 400 BC, which performs a simple and yet practical transposition cipher, Caesar cipher from 50 BC, which performs a simple letter substitution method by replacing each letter in the message with the one appearing alphabetically three places apart, polyalphabetic ciphers in the 15th century, which are based on superpositions of Caesar ciphers with different shifts, and the famous Enigma machine of Arthur Scherbius developed after World War I, which uses electromechanical rotor technology to generate an encrypting sequence with an extremely long period of substitutions. Long story short, history is full of battles between cryptographers (codemakers) and cryptanalysts (codebreakers), with either side taking the lead at different points in time [14]. Considering the rise of physics-based and unconditionally secure key distribution systems, one can ask: Is it game over for codebreakers? After this quick introduction, this section presents the key distribution and keyless encryption approaches and their practical challenges. Finally, we briefly discuss the potential of PHY security approaches.
@@ IV-A Key Distribution Problem
On the search for truly unbreakable ciphers, the one-time pad (or the Vernam cipher [15]), illustrated in Fig. 3, shines out as a potential candidate to secure communications, once and for all. It was developed in the early 20th century and mathematically proven unbreakable by Shannon in the late 1940s [16]. However, to ensure an unbreakable cipher using the one-time pad, the encryption key must be as long as the plaintext to be encrypted, it must be completely random, it must never be reused in whole or in part, and it must be kept completely secret between communicating parties. In other words, according to the terminology of Shannon, the entropy of the key ($K$) must be higher than that of the message ($M$). This brings out the secure key distribution problem. The primary issue in secure communication is the key distribution problem [14, 17]. A naive solution would be distributing the keys physically, as done by certain banks in the 1970s using trusted employees to convey decryption keys to their customers; however, this is not technically and logistically feasible by today’s standards. No matter how secure a cipher is in theory, it is challenged by key distribution problems in practice. In other words, if two parties want to communicate securely, they must rely on a third party to deliver the key, which is the Achilles heel of the system. Interestingly, the entire key distribution problem embodies a catch-22 situation, a paradoxical scenario one cannot escape due to its contradictory rules or limitations. When two individuals aim to exchange a secret message, the sender must first encrypt it. However, for encrypting the secret message, the sender must utilize a key that remains a secret. This dilemma presents the challenge of transmitting the secret key to the receiver to enable the transmission of the encrypted secret message. Consequently, to effectively communicate a secret message, the two users must already share a secret (the key). Despite the most revolutionary development in 20th-century cryptography being the creation of effective mathematical techniques to overcome this problem, the extensive implementation of quantum computers in the future might challenge their computational (conditional) security. At this juncture, a pertinent question arises: Can a secret message be securely communicated without the need for exchanging a key?
@@ IV-B Keyless Encryption and Practical Issues
For 2000 years, the distribution of keys has been regarded as an indisputable aspect of secure communication. However, challenging this dogma, the doubly encrypted scheme (Shamir’s three-pass or no-key protocol developed in the 1970s) enables the exchange of secure messages without the need to exchange a key [18]. For the first time, we think that key exchange might not be an inevitable part of secure communication! The three-pass protocol can be better understood by invoking the well-known mechanical analogy of a mail carrier box with two padlocks, as illustrated in Fig. 4 [19, 20]. Assume that Alice intends to send a highly confidential message to Bob, accounting for an unreliable carrier service. She places the message in a box, secures it with her green padlock (operation $A$), and then dispatches it to Bob (Stage 1). Upon receiving the box, Bob cannot unveil the secret message as he lacks the key to Alice’s padlock. Conversely, he adds his red padlock (operation $B$) and returns the double padlocked box to Alice (Stage 2). At this point, Alice removes her green padlock and forwards the box back to Bob (Stage 3). The returned box only contains Bob’s padlock, which he removes to access the secret message. Thus, secure transmission has been achieved in three stages, all without a key involvement! It appears that the problem of key distribution is resolved, yet several fundamental challenges remain to be tackled. These challenges encompass the development of practical mathematical tools (involving classical physical operators) that incorporate randomness and adhere to the specific sequence of Alice and Bob’s locking/unlocking operations, ensuring an unconditionally secure and implementable three-pass scheme. In this context, it is necessary to ensure that even if Eve, who represents one of the central components of this unreliable carrier service between Alice and Bob, observes all three publicly shared messages, she cannot recover the concealed message. Additionally, it is crucial that Alice and Bob share no information about their actual operators $A$ and $B$. The practical problems associated with the three-pass protocol have resulted in the development of the Diffie-Hellman-Merkle (DHM) key exchange scheme in 1976. This scheme involves modular arithmetic $(Y^{x}(\mathrm{mod}\,P))$ to facilitate key exchange between Alice and Bob through public discussion [14]. Nevertheless, the three-pass protocol and the DHM key exchange scheme share a significant drawback: Alice and Bob must exchange messages and wait for each other’s responses. The invention of public key cryptography (also known as asymmetric cryptography) by Rivest, Shamir, and Adleman in 1977 not only addressed this waiting problem stemming from public discussions but also overcame the challenge of key distribution by leveraging public and private keys, which are based on huge prime numbers. To illustrate with a mechanical analogy, Alice creates numerous replicas of her padlock and distributes them to various mail offices in this scenario. A sender wishing to convey a secret message to Alice places it in a box and employs a replica of Alice’s padlock (public key) before sending it. Only Alice can unlock this box, as she is the sole possessor of the private key for these padlocks. In essence, anyone can secure a padlock by simply snapping it shut; however, only the person with the private key can open it. A notable drawback of asymmetric encryption is its demand for more time and computing power, which may not be suited for devices with limited memory and computational resources. On the other hand, asymmetric encryption can serve as a secure method for exchanging a symmetric key to be utilized in bulk encryption on suitable devices [21].

### v3

@@ III-A Fundamentals of KLJN
One can ask the following question regarding the KLJN system: Compared to the bulletproof security provided by quantum weirdness, can a simple wire that is freely exposed to the public, with measurable voltage and current waveforms, offer unconditional security for eavesdroppers with unlimited measurement capabilities? The answer is yes, thanks to the laws of physics (thermodynamics). In other words, like its quantum-based counterpart, violating the security of KLJN requires violating specific laws of physics.
One of the reasons why our community has not fully understood the scheme of KLJN in the past 15 years is due to its multidisciplinary nature, which includes statistical physics, circuit noise theory, stochastic signals, and information-theoretic security. In the following subsection, we will discuss the basic technical operation principles of the KLJN scheme.
@@ III-B KLJN Operation
which can alternatively be expressed in units of dBm/Hz or watts/Hz considering a reference impedance. In (1), $R_{A}$ and $R_{B}$ represent the randomly selected resistors by Alice and Bob, respectively, that is, $R_{A},R_{B}\in\left\{R_{L},R_{H}\right\}$. Accordingly, for the four possible cases of selected Alice/Bob bits, namely $00,01,10,$ and $11$, where the first and second bits respectively stand for the selected bits by Alice and Bob, voltage waveform samples on the line will be Gaussian distributed random variables with different variances. The corresponding variance values of the noise samples are given as follows
@@
$\sigma_{11}^{2}$ $=4kT\frac{R_{H}R_{H}}{R_{H}+R_{H}}\Delta f=4kTR_{L}\frac{\alpha}{2}\Delta f,$ (2)
where $\alpha$ is the ratio of resistances, that is, $R_{H}=\alpha R_{L}$. As illustrated in Fig. 1, the fluctuations on the line will depend on the selected resistance values at both sides of the system. Specifically, the fluctuations on the line will be high if both Alice and Both select the large resistance $R_{H}$. On the other hand, if both Alice and Bob select the small resistance $R_{L}$, the fluctuations will be small. What about the intermediate level? If one user (Alice or Bob) selects the large resistance while the other selects the small one, the noise variance on the line takes an intermediate value, which lies at the heart of the KLJN operation [10]. This intermediate noise variance value creates an indeterminacy for an eavesdropper (Eve), thus paving the way for unconditional security. Before providing more detailed discussions on this important topic, let us give a numerical example of the noise variances of different cases. Specifically, for the value of $\alpha=10$, noise variances on the line for the cases of intermediate and high fluctuations would be $1.8182$ and $10$ times greater than that of small fluctuations, respectively.
An unconditionally secure bit exchange occurs when the bit values (or equivalently resistances) at the two ends of the KLJN system differ [^1] Although a random key is established between Alice and Bob by the KLJN system, it has been widely regarded as a key exchanger in literature due to its unique architecture.. For this case, an intermediate level is observed for the noise voltage variance (mean-square noise voltage), as shown in Fig. 1. Even though this intermediate level can be easily detected by Eve from a sufficient number of noise samples taken over the line, the specific contributions of Alice and Bob cannot be understood. This confusion on Alice’s and Bob’s resistances, created deliberately for Eve, ensures unconditional security at the level of quantum-based key distribution systems. In other words, the KLJN scheme consistently creates an indeterminacy for external observers, rendering the intermediate noise variance meaningless. In contrast, the cases of $00$ and $11$, respectively, provide the lowest and highest noise variances on the link. In this case, an eavesdropper can identify Alice/Bob bits from its measurements without a significant effort, corresponding to a non-secure bit exchange.
As stated earlier, under ideal conditions, the cases of $01$ and $10$ produce equivalent noise variances for the noise samples. As a result, the specific locations of Alice/Bob bits cannot be determined from measurements taken on the line between the users. This unique property of KLJN provides a sort of powerful encryption and ensures absolute (unbreakable) security. For Alice and Bob, determining their partner’s selected bit is straightforward for all four cases using voltage and/or current samples and the knowledge of their own bit. However, they have to take a sufficient number of noise samples for each transmission interval to reduce their bit errors. While we provide brief coverage of this issue in the following subsection, interested readers are referred to [3] for a comprehensive theoretical evaluation of KLJN’s bit error probability (BEP) and its different types of detectors exploiting noise voltage and/or current samples.
It is worth noting that the switching mechanism of the KLJN scheme among low and high resistances has conceptual similarities with the index modulation (IM) method [11], which embeds information in certain system entities such as antennas, subcarriers, and time slots. To be more specific, a KLJN communicator performs a unique form of IM by considering the indices of its available two resistors, or equivalently, for three possible noise voltage (or current) power spectral densities; however, it requires a no-wave solution for security. As in all IM systems, index selection does not consume power, and the primary task of the users is to extract the information embedded in the corresponding system entity through IM, which is the noise variance for the case of the KLJN scheme. In this context, selecting the number of noise samples per bit is critical, as discussed in the following subsection.
While concluding this subsection here, Section VI discusses the state-of-the-art KLJN solutions and their security vulnerabilities.
@@ III-C Conquering the Noise: KLJN Detection
Noise: an enemy to be dealt with for achieving reliable and secure communication. Deviating from mainstream engineering, we can ask the following question: is it possible to use the noise itself, for instance, as biology does using noise-alike signals as the information carrier, for communication purposes? KLJN uses the noise itself for this purpose; as a result, we need to explore practical ways to extract the information embedded in this noise. In this subsection, we shed light on this issue.
To determine their partner’s bit in the KLJN system, Alice and Bob take samples from the thermal noise voltage (or current) on the wire for each bit duration $(T_{b})$. The total number of “uncorrelated” noise samples taken per bit, denoted by $N$, can be derived from the Wiener–Khinchin theorem, considering the sinc-type time autocorrelation function of the band-limited white noise. Accordingly, for a noise bandwidth of $\Delta f$ Hz, a maximum of $N=2T_{b}\Delta f$ noise samples can be taken for each bit duration to ensure that the samples are statistically independent, which is also consistent with the Nyquist’s sampling theorem. In other words, in accordance with the sampling theorem, faster-than-Nyquist sampling would result in correlated noise samples. Consequently, to increase $N$ to improve the accuracy of noise variance estimates, one needs to increase $T_{b}$; however, this reduces the bit rate given by $R_{b}=1/T_{b}$. This provides an interesting trade-off between the bit error rate (BER) and data rate for the KLJN system. Another attractive feature of KLJN is that no matter how strong Eve’s signal processing capabilities are, the same limit also applies to her; she cannot gain further knowledge by expanding her observation interval or increasing the number of noise samples.
@@
$\hat{\sigma}^{2}=\frac{1}{N}\sum\nolimits_{k=1}^{N}x_{k}^{2}.$ (3)
An effective way for Alice and Bob to detect their partner’s bit is to employ a threshold-based detector, as shown in Fig. 2, by employing two threshold values, $\gamma_{1}$ and $\gamma_{2}$. At this point, considering the value of $\hat{\sigma}^{2}$ as well as their own bit, Alice and Bob can determine their partner’s bit. For instance, assume that Alice’s bit is 00. In this case, she considers the first threshold, $\gamma_{1}$, to detect Bob’s bit. If $\hat{\sigma}^{2}<\gamma$, her decision becomes 00, and she marks the corresponding bit as an insecure one since she assumes that the selected bit combination is $00$. Particularly, she is aware of the fact that a third party can easily identify the case of $00$ through measurements on the line.
On the contrary, if $\hat{\sigma}^{2}>\gamma$, her decision becomes $1$, which corresponds to an unconditionally secure bit exchange (generation). Since Alice and Bob’s bits are different for this secure case, one can flip his/her bits to create a shared and secure key after the transmission. It is worth noting that for the detection scheme of Fig. 2, the error events between $00$ and $01/10$ dominate the overall BEP due to the proximity of the corresponding variance values on the decision space. Employing joint or alternating noise current samples, which provide an opposite relationship between noise variances compared to those given by voltage samples, this problem can be alleviated. As a result, much lower BEP can be achieved using more sophisticated detectors.
It can be shown that the BEP of the KLJN scheme exponentially decays with the number of noise samples per bit, that is, $P_{b}\propto\exp(-N)$, which is promising to obtain very low BEP values for secure key generation between two partners. Furthermore, lowering the BEP does not compromise the unconditional security of the system. A detailed theoretical framework is represented in [3] by considering decision errors due to the randomness of the estimated noise variance. Furthermore, two new detectors are also presented in this study by exploiting joint voltage and noise samples further to improve the BEP performance of the KLJN scheme.
@@ IV In Pursuit of Unconditional Security
One of the objectives during the introduction of KLJN in the early 2000s had been to challenge the following status quo on unconditional security: Only quantum physics can provide unconditional security for the key distribution. At this point, the conceptual difference between post-quantum cryptography and quantum cryptography is worth noting. Specifically, post-quantum cryptography, also known as quantum-resistant or quantum-safe cryptography, deals with quantum-resistant cryptographic algorithms, which are implemented on existing platforms and derive their security through mathematical complexity. In principle, these algorithms will be resilient to attacks from quantum computers thanks to their highly complicated mathematical frameworks [12]. On the other hand, quantum cryptography, specifically quantum key distribution (QKD), uses principles of quantum mechanics to distribute the keys to be used later in a suitable form of cryptography and is genuinely unbreakable, unlike mathematical encryption. In other words, the security level provided by QKD is beyond the attackers of the future who possess quantum computers and even have infinite computational power [^2] This strong assertion holds when we employ one-time pad ciphers with keys established using QKD, which will be discussed subsequently.. In what follows, we dive into a historical journey on secure communication to catch valuable insights on the motivation of these physics-based key generation systems, KLJN and QKD, and their limitations in practice.
The desire to communicate secretly has existed as long as writing and encrypted communication methods were developed in many ancient societies [13]. Notable historical examples include the scytale of Spartans from 400 BC, which performs a simple and yet practical transposition cipher, Caesar cipher from 50 BC, which performs a simple letter substitution method by replacing each letter in the message with the one appearing alphabetically three places apart, polyalphabetic ciphers in the 15th century, which are based on superpositions of Caesar ciphers with different shifts, and the famous Enigma machine of Arthur Scherbius developed after World War I, which uses electromechanical rotor technology to generate an encrypting sequence with an extremely long period of substitutions. Long story short, history is full of battles between cryptographers (codemakers) and cryptanalysts (codebreakers), with either side taking the lead at different points in time [14]. Considering the rise of physics-based and unconditionally secure key distribution systems, one can ask: Is it game over for codebreakers?
After this quick introduction, this section presents the key distribution and keyless encryption approaches and their practical challenges. Finally, we briefly discuss the potential of PHY security approaches.
@@ IV-A Key Distribution Problem
On the search for truly unbreakable ciphers, the one-time pad (or the Vernam cipher [15]), illustrated in Fig. 3, shines out as a potential candidate to secure communications, once and for all. It was developed in the early 20th century and mathematically proven unbreakable by Shannon in the late 1940s [16]. However, to ensure an unbreakable cipher using the one-time pad, the encryption key must be as long as the plaintext to be encrypted, it must be completely random, it must never be reused in whole or in part, and it must be kept completely secret between communicating parties. In other words, according to the terminology of Shannon, the entropy of the key ($K$) must be higher than that of the message ($M$). This brings out the secure key distribution problem.
The primary issue in secure communication is the key distribution problem [14, 17]. A naive solution would be distributing the keys physically, as done by certain banks in the 1970s using trusted employees to convey decryption keys to their customers; however, this is not technically and logistically feasible by today’s standards. No matter how secure a cipher is in theory, it is challenged by key distribution problems in practice. In other words, if two parties want to communicate securely, they must rely on a third party to deliver the key, which is the Achilles heel of the system.
Interestingly, the entire key distribution problem embodies a catch-22 situation, a paradoxical scenario one cannot escape due to its contradictory rules or limitations. When two individuals aim to exchange a secret message, the sender must first encrypt it. However, for encrypting the secret message, the sender must utilize a key that remains a secret. This dilemma presents the challenge of transmitting the secret key to the receiver to enable the transmission of the encrypted secret message. Consequently, to effectively communicate a secret message, the two users must already share a secret (the key).
Despite the most revolutionary development in 20th-century cryptography being the creation of effective mathematical techniques to overcome this problem, the extensive implementation of quantum computers in the future might challenge their computational (conditional) security. At this juncture, a pertinent question arises: Can a secret message be securely communicated without the need for exchanging a key?
@@ IV-B Keyless Encryption and Practical Issues
For 2000 years, the distribution of keys has been regarded as an indisputable aspect of secure communication. However, challenging this dogma, the doubly encrypted scheme (Shamir’s three-pass or no-key protocol developed in the 1970s) enables the exchange of secure messages without the need to exchange a key [18]. For the first time, we think that key exchange might not be an inevitable part of secure communication! The three-pass protocol can be better understood by invoking the well-known mechanical analogy of a mail carrier box with two padlocks, as illustrated in Fig. 4 [19, 20]. Assume that Alice intends to send a highly confidential message to Bob, accounting for an unreliable carrier service. She places the message in a box, secures it with her green padlock (operation $A$), and then dispatches it to Bob (Stage 1). Upon receiving the box, Bob cannot unveil the secret message as he lacks the key to Alice’s padlock. Conversely, he adds his red padlock (operation $B$) and returns the double padlocked box to Alice (Stage 2). At this point, Alice removes her green padlock and forwards the box back to Bob (Stage 3). The returned box only contains Bob’s padlock, which he removes to access the secret message. Thus, secure transmission has been achieved in three stages, all without a key involvement!
It appears that the problem of key distribution is resolved, yet several fundamental challenges remain to be tackled. These challenges encompass the development of practical mathematical tools (involving classical physical operators) that incorporate randomness and adhere to the specific sequence of Alice and Bob’s locking/unlocking operations, ensuring an unconditionally secure and implementable three-pass scheme. In this context, it is necessary to ensure that even if Eve, who represents one of the central components of this unreliable carrier service between Alice and Bob, observes all three publicly shared messages, she cannot recover the concealed message. Additionally, it is crucial that Alice and Bob share no information about their actual operators $A$ and $B$.
@@
The practical problems associated with the three-pass protocol have resulted in the development of the Diffie-Hellman-Merkle (DHM) key exchange scheme in 1976. This scheme involves modular arithmetic $(Y^{x}(\mathrm{mod}\,P))$ to facilitate key exchange between Alice and Bob through public discussion [14]. Nevertheless, the three-pass protocol and the DHM key exchange scheme share a significant drawback: Alice and Bob must exchange messages and wait for each other’s responses. The invention of public key cryptography (also known as asymmetric cryptography) by Rivest, Shamir, and Adleman in 1977 not only addressed this waiting problem stemming from public discussions but also overcame the challenge of key distribution by leveraging public and private keys, which are based on huge prime numbers. To illustrate with a mechanical analogy, Alice creates numerous replicas of her padlock and distributes them to various mail offices in this scenario. A sender wishing to convey a secret message to Alice places it in a box and employs a replica of Alice’s padlock (public key) before sending it. Only Alice can unlock this box, as she is the sole possessor of the private key for these padlocks. In essence, anyone can secure a padlock by simply snapping it shut; however, only the person with the private key can open it. A notable drawback of asymmetric encryption is its demand for more time and computing power, which may not be suited for devices with limited memory and computational resources. On the other hand, asymmetric encryption can serve as a secure method for exchanging a symmetric key to be utilized in bulk encryption on suitable devices [21].